OpenStudy (anonymous):

Medal to who ever helps Part 1: Solve the following system of equations. Please show your work to receive full credit. x - y = 10 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice.

OpenStudy (anonymous):

I'LL CHOOSE THE ELIMINATION METHOD. LEAST AMOUNT OF WORK. (11 STEPS) x - y = 10 2x + y = 2 ADD ------------- 3X=12 X=12/3 X=4 ANS. 4-Y=10 -Y=10-4 -Y=6 Y=-6 ANS. PROOF: 2*4+(-6)=2 8-6=2 2=2

OpenStudy (anonymous):

Thank u

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

Select the system of equations that corresponds to the given graph. A coordinate plane shows two lines on the graph. One line crosses the y axis at (0, 2) and passes through the point (2,1). The other line crosses the y axis at (0,-2) and passes through the point (2,-1). -2x - y = 4 2x + y = -2 -8x + 4y = -16 -4x + 2y = -8 x + 2y = 4 2x - 4y = 8 2x + 3y = 6 3x + 2y = 4

OpenStudy (anonymous):

Help